FINITELY SMOOTH REINHARDT DOMAINS WITH NON-COMPACT AUTOMORPHISM GROUP

STEVEN G. KRANTZ ALEXANDER V. ISAEV

Abstract. We give a complete description of bounded Reinhardt domains of finite boundary smoothness that have non-compact automorphism group. As part of this program, we show that the classification of domains with non-compact automorphism group and having only finite boundary smoothness is considerably more complicated than the classification of such domains that have infinitely smooth boundary.